Mr. Johnson has 11 quarts of punch that is 16% lemonade. He wants to makee some punch containing 52% lemonade. How much lemonade should he add?
Answer by scott8148(6628)

let x="qts lemonade added" __ 11(16%)+x(100%)=(11+x)(52%) __ multiplying by 100 to clear percentages

176+100x=572+52x __ 48x=396